What is the probability of getting 8 heads in a row? Assume that each coin toss is independent of the other.
Answer #1

as probability of head on a trail =1/2

therefore from independence

P( getting 8 heads in a row )=P(head)*P(head)*..P(head) =(1/2)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2)

=1/256